OpenStudy (anonymous):

What is the y-intercept of the line? y=-2x-8

OpenStudy (secret-ninja):

The equation for graphing is y=mx+b and the "+b" is the y intercept. whether it is positive or negative, makes it goes upward or downward.

OpenStudy (secret-ninja):

What do you think is the y intercept?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

The y-intercept is where the line crosses the y-axis, meaning what is the value of y when x =0?

OpenStudy (secret-ninja):

You don't have to solve anything, it is already in the equation.  See?

OpenStudy (secret-ninja):

So what is your y intercept?

OpenStudy (secret-ninja):



Directrix (directrix):

To get on the y-axis, the x value has to be 0. Take the equation y=-2x-8 and substitute 0 for x. y = -2*0 - 8 = 0 - 8 = ? @_Invalid._.user

OpenStudy (anonymous):

8

Directrix (directrix):

What is 0 - 8 = ?

OpenStudy (anonymous):

-8

Directrix (directrix):

Yes. The y-intercept is -8. The point coordinates can be written as (0, -8)

OpenStudy (secret-ninja):

Good job! :D

OpenStudy (anonymous):

whats the slope, slope= and whats the why intersect